OpenStudy (cookiemonster2412):

Find the equation of the line given the following information. Slope and y-intercept. m=3 point(0,7) m=2/5 point(0,1)

OpenStudy (cookiemonster2412):

@One098

OpenStudy (cookiemonster2412):

@TheRaggedyDoctor

OpenStudy (theraggedydoctor):

Wait... why's there two slopes? xD

OpenStudy (cookiemonster2412):

its two diff problems

OpenStudy (theraggedydoctor):

Forget I said that. I was thinking of something else XD

OpenStudy (theraggedydoctor):

But since you already know the slope, you can just plug in the coordinates for x and y to find the y intercept (b). Sooooo: 7 = 3(0) + b 7 = 0 + b 7 = b So the first one is: y = 3x + 7 1 = 2/5(0) + b 1 = 0 + b 1 = b The second problem would be: y = 2/5x + 1

OpenStudy (cookiemonster2412):

oooo u smart spap

OpenStudy (theraggedydoctor):

I'm smart in math, that's about it xD

OpenStudy (cookiemonster2412):

i doubt that o.o

OpenStudy (theraggedydoctor):

It's true xD and theology perhaps

OpenStudy (cookiemonster2412):

thats not wat ive heard ;) xDD